OpenStudy (anonymous):

tell me factors of 456 plz

OpenStudy (anonymous):

A factor is anything that 456 can be divided by.

OpenStudy (anonymous):

how do you do factors?

OpenStudy (anonymous):

\[ 456 = 2^3* 3 * 19 \]

OpenStudy (cherylim23):

You factor integers using prime numbers. Try the first prime number, 2, and divide 456 by it. You should get 228. Keep dividing by 2 until you get an odd number. 228/2 = 114, 114/2 = 57. Now that you can't divide by 2, try the next prime number, 3. 57/3 = 19 At this stage, the dividend, 19, itself is a prime number so you're nearly done. How many times have you divided by 2? 3 times. So your answer should be 2^3 x 3 x 19 That's how Ron.mystery got his answer

OpenStudy (anonymous):

@Cherylim23 You explained that really well. :)

OpenStudy (cherylim23):

Thanks, maths begins all with foundation.

OpenStudy (anonymous):

\[\huge 456 = 2^3 * 3* 19 \]

OpenStudy (anonymous):

are the factors

ganeshie8 (ganeshie8):

456 = \(\color{purple}{1}\) x 456 R 0 456 = \(\color{purple}{2}\) x 228 R 0 456 = \(\color{purple}{3}\) x 152 R 0 456 = \(\color{purple}{4}\) x 114 R 0 456 = \(\color{purple}{6}\) x 76 R 0 456 = \(\color{purple}{8}\) x 57 R 0 456 = \(\color{purple}{12}\) x 38 R 0 456 = \(\color{purple}{19}\) x 24 R 0 456 = \(\color{purple}{24}\) x 19 R 0 456 = \(\color{purple}{38}\) x 12 R 0 456 = \(\color{purple}{57}\) x 8 R 0 456 = \(\color{purple}{76}\) x 6 R 0 456 = \(\color{purple}{114}\) x 4 R 0 456 = \(\color{purple}{152}\) x 3 R 0 456 = \(\color{purple}{228}\) x 2 R 0 456 = \(\color{purple}{456}\) x 1 R 0

OpenStudy (anonymous):

haha Boss.

ganeshie8 (ganeshie8):

just playing with latex... hehe btw purple ones are the factors of 456

OpenStudy (anonymous):

Yeah, I know. Great job! haha It's fun :D

OpenStudy (anonymous):

tnx guys